A commercial bank has reserves of $64, loans of $521 and checkable deposits of $585. The bank experiences a cash outflow of $11. If the required reserve ratio is 7%, what are the bank's excess reserves after the outflow?
Answer #1

Answer

New checkable deposits =585-11

=574

Required reserves =checkable deposits * required reserve ratio

=574*0.07

=40.18

New reserves =old reserves - cash outflow

=64-11

=53

Excess reserves =new reserves - required reserves

=53-40.18

=12.82

The excess reserves are $12.82